hi I do not undrestand solving using substitution method for example
x=y-1 x=y+6 r=-3s
y=4-2x y=-2-x r=10-4s









Answer by checkley71(8405)

x=y-1 now substitute (y-1) for x in the other equation & solve for y:
y=4-2(y-1)
y=4-2y+1
y+2y=5
3y=5
y=5/3 answer.
now replace y with 5/3 & solve for x:
x=5/3-1
x=5/3-3/3
x=2/3 answer.
-------------------------------------------------
x=y+6 now replace x in the othere equation & solve for y:
y=2-(y+6)
y=2-y-6
y+y=-4
2y=-4
y=-4/2
y=-2 answer.
x=-2+6
x=4 answer.
--------------------------------------------------------
Now do the same with the next set of equations: